A round pizza is baking in the oven. As the pizza gets warmer, its radius increases at a rate of 0.01 centimeters per minute. When the diameter of the pizza is 50 centimeters, what is the rate that the pizza's area is increasing?
Your solution myst include a sketch of the situation and labeling all given quantities. Leave your final answer in terms of pi.
1
Expert's answer
2020-06-10T17:31:17-0400

The area of the round pizza is "A=\pi r^2" , where "r" is he radius.


When the diameter is "50" cm, the radius of the pizza is "r=\frac{50}{2}=25" cm.


Differentiate "A" with respect to "t" as,


"\frac{dA}{dt}=\frac{d}{dt}(\pi r^2)"


"=\pi (2r)\frac{dr}{dt}"


"=2\pi r\frac{dr}{dt}"


Plug "r=25" and "\frac{dr}{dt}=0.01" to find the rate of increase of area as,


"\frac{dA}{dt}=2\pi(25)(0.01)"


"=50\pi(0.01)"


"=\frac{1}{2}\pi"


Therefore, the rate of increase of area is "\frac{dA}{dt}=\frac{1}{2}\pi=0.5\pi" "cm^2\/min"